i need help with my equations in reference to if my answers are correct
(1) 7x+3=5x+17
(2) -11-3c=4+2c
(3) 3-2a=12-9
(4) 2(3x+7)=16
(5) 3x=5(9-x)
(6) 8-5(2-x)=-17

Answer by stanbon(57290)

Comment: You should post your answers.
It saves us time.
-----------------------
(1) 7x+3=5x+17
x = 7
--------------
(2) -11-3c=4+2c
x = -3
--------------
(3) 3-2a=12-9
a = 0
--------------
(4) 2(3x+7)=16
x = 1/3
----------------
(5) 3x=5(9-x)
x = 45/8
===============
(6) 8-5(2-x)=-17
x = -3
=====================
Cheers,
Stan H.
===============